### v1

# Do Schwartz Higher-Order Values Help Sentence-Level Human Value Detection? When Hard Gating Hurts

Abstract
Sentence-level human value detection is typically framed as multi-label classification over Schwartz values, but it remains unclear whether Schwartz higher-order (HO) categories provide usable structure. We study this under a strict compute-frugal budget (single 8 GB GPU) on ValueEval’24 / ValuesML (74K English sentences). We compare (i) direct supervised transformers, (ii) HO values pipelines that enforce the hierarchy with hard masks, and (iii) Presence HO values cascades, alongside low-cost add-ons (lexica, short context, topics), label-wise threshold tuning, small instruction-tuned LLM baselines (10B), QLoRA, and simple ensembles. HO categories are learnable from single sentences (e.g., the easiest bipolar pair reaches Macro-), but hard hierarchical gating is not a reliable win: it often reduces end-task Macro- via error compounding and recall suppression. In contrast, label-wise threshold tuning is a high-leverage knob (up to Macro-), and small transformer ensembles provide the most consistent additional gains (up to Macro-). Small LLMs lag behind supervised encoders as stand-alone systems, yet can contribute complementary errors in cross-family ensembles. Overall, HO structure is useful descriptively, but enforcing it with hard gates hurts sentence-level value detection; robust improvements come from calibration and lightweight ensembling.
keywords:
Morality detection , Human values , Schwartz value theory , Sentence-level classification , Transformer models , Ensemble learning , Large language models[upv]organization=PRHLT Research Center, Universitat Politècnica de València, city=Valencia, postcode=46022, country=Spain
[uev]organization=School of Science, Engineering and Design, Universidad Europea de Valencia, city=Valencia, postcode=46010, country=Spain
[valgrai]organization=Valencian Graduate School and Research Network of Artificial Intelligence (ValgrAI)
1 Introduction
Human values are enduring guiding principles that shape what people consider important, desirable, or worth protecting (Rokeach, 1973; Schwartz, 1992; Bardi and Schwartz, 2003). Because values are often expressed implicitly in language, detecting them in text matters for computational social science and NLP tasks that analyze public discourse, persuasion, stance, framing, and argumentation at scale (Lazer et al., 2009). Recent surveys of computational morality and value modeling summarize key approaches (lexicon-based signals, supervised classifiers, LLM-centric methods) and emphasize persistent challenges such as contextual ambiguity and domain sensitivity (Reinig et al., 2024). Related work on moral language in political and social media discourse shows that moral/value cues are typically sparse, indirect, and context-dependent (Haidt and Joseph, 2004; Johnson and Goldwasser, 2018).
Among social-science value frameworks, Schwartz’s theory is widely adopted and empirically validated (Schwartz, 1992). The refined theory defines 19 basic values and groups them into higher-order (HO) categories (e.g., Openness to Change vs. Conservation) that capture compatibilities and conflicts (Schwartz, 2012). This hierarchy suggests a potential inductive bias for predicting fine-grained values when labels are sparse or ambiguous. Figure 1 shows the circular motivational continuum of the 19 values, where adjacent values are compatible and opposing values are in conflict (Schwartz et al., 2012).
Recent shared tasks operationalize value detection as sentence-level, multi-label prediction, enabling controlled comparisons (Kiesel et al., 2023). The Touché 2024 Human Value Detection task (ValueEval’24) is built on ValuesML, where sentences are annotated for which Schwartz values they express and whether a value is attained or constrained (The ValuesML Team, 2024; Touché, 2024). Recent work highlights machine learning challenges and benchmark-driven evaluation, showing how dataset design and label distributions shape performance (Rink et al., 2025). These benchmarks reveal realistic difficulties: a sentence may express none, one, or many values; evidence is often implicit and lexically diffuse; and label prevalence is highly imbalanced (Kiesel et al., 2023). These properties strain standard multi-label pipelines and make calibration decisions (thresholds, probability reliability) especially important (Tsoumakas and Katakis, 2007; Guo et al., 2017; Silva Filho et al., 2023).
A natural question is whether Schwartz’s HO structure improves fine-grained sentence-level detection. Hierarchical classification can help by injecting structure, constraining hypotheses, and sharing statistical strength (Silla and Freitas, 2011). But in noisy sentence-level settings, hard constraints can amplify upstream errors: if a parent prediction is uncertain, strict gating can suppress true positives and reduce recall on already sparse labels. This tension is tied to calibration, since hierarchical pipelines are sensitive to thresholds and probability miscalibration (Valmadre, 2022).
We address this tension by evaluating when HO categories help and how to integrate them. Under a compute-frugal setting, we compare: (i) direct multi-label prediction, (ii) a CategoryValues hierarchy that constrains fine-grained outputs, and (iii) a PresenceCategoryValues cascade that first filters sentences predicted to contain any value. We also test low-cost levers that often matter in practice, including threshold calibration and simple ensembling (Wolpert, 1992; Breiman, 1996; Freund and Schapire, 1997; Breiman, 2001). Finally, we benchmark compact instruction-tuned LLMs under the same budget, motivated by evidence that prompting and instruction tuning can be competitive without task-specific architectural changes (Brown et al., 2020; Ouyang et al., 2022; Chung et al., 2024).
Research questions
We structure the study around the following research questions:
RQ1.
Are HO values learnable from single sentences? Can we reliably detect the eight Schwartz HO categories from a single sentence, and which compute-frugal signals and model families work best?
RQ2.
Do HO gates help downstream basic-value prediction? Does inserting an HO category detector as a gate before predicting the 19 basic values improve out-of-sample Macro- compared to a single-stage Direct model?
RQ3.
Does a PresenceCategory cascade improve over Category-only? If we add a Presence gate before the HO gate (Presence CategoryValues), does this hierarchy outperform (a) Direct prediction and (b) CategoryValues on the test set?
RQ4.
What low-cost knobs actually move the needle? Across HO detection and the hierarchical pipeline, which lightweight signals (lexica, topic vectors, short local context) and which calibration/ensembling choices yield statistically supported gains under fixed compute?
RQ5.
Where do small LLMs fit? Under the same budget, how do instruction-tuned 10B LLMs (zero-shot, few-shot, and QLoRA) compare to supervised DeBERTa-based models for HO detection and for the hierarchy-driven pipeline?
Our contributions are:
A focused study of whether Schwartz HO categories improve sentence-level value detection on ValueEval’24/ValuesML, including analyses by canonical bipolar HO pairs.
A comparison of HO-aware strategies (conditioning, hard gating, cascades) that quantifies when hierarchy helps and when it fails due to error propagation.
A compute-frugal evaluation of practical improvements—threshold tuning, calibration-aware decisions, and small ensembles—that yield gains without increasing model size or training cost.
Overall, we provide a practical blueprint for using value structure in NLP pipelines, and show that hard hierarchical constraints are brittle in this sentence-level setting, while calibration and small ensembles deliver the most reliable gains under a compute-frugal budget.
The rest of the paper is organized as follows. Section 2 reviews prior work on value and moral language detection, benchmarks, hierarchical and multi-label learning, calibration/ensembling, and the use of transformers and instruction-tuned LLMs. Section 3 describes the task, dataset, model variants, and compute-frugal protocol. Section 4 reports results and analyzes when HO structure helps (or hurts). Section 5 discusses implications and answers the research questions. Section 6 concludes and outlines future work. Tables and figures labeled Sx (e.g., Table S4) refer to the Supplementary Material.
2 Related work
2.1 Human values and moral frameworks in NLP
Human values are commonly operationalized as relatively stable guiding principles that shape preferences and judgments (Rokeach, 1973; Schwartz, 1992; Bardi and Schwartz, 2003). For an NLP-oriented synthesis of how morality/value constructs are used in text analysis, see Reinig et al. (2024). Among taxonomies, Schwartz’s theory is especially attractive for NLP because it provides (i) a refined, fine-grained set of basic values and (ii) a principled HO organization that captures compatibilities and conflicts (Schwartz et al., 2012). Computational work on moral language often draws on Moral Foundations Theory (MFT) to study moral rhetoric in political and social discourse (Graham et al., 2009). Although MFT and Schwartz address different constructs, both highlight a key modeling challenge: moral/value signals in text are often indirect, sparse, and diffuse rather than explicitly labeled (Graham et al., 2011; Haidt, 2012). Beyond classification, NLP has also been used to elicit structured value representations for downstream systems, such as extracting value promotion schemes (García-Rodríguez et al., 2025).
An extensive line of work explores feature- and lexicon-driven prediction as an interpretable, low-cost alternative to heavy end-to-end models (Hopp et al., 2021; Hoover et al., 2020). Araque et al. (2020) propose MoralStrength, which extends the Moral Foundations Dictionary with embedding-based similarity. González-Santos et al. (2023) study moral foundations assignment in the movie domain using word embeddings and semantic similarity. These studies motivate our compute-frugal perspective: lightweight signals can be useful, but performance depends on how prior structure is injected.
Recent work expands the scope of value detection beyond classic lexicon settings to online community discourse and multimodal platforms, and proposes methods for large text collections. For example, value expressions are analyzed in online communities (Borenstein et al., 2025) and in multimodal influencer content (Starovolsky-Shitrit et al., 2025), while context-dependent markup schemes are proposed for large-scale value/sentiment detection in social media corpora (Rink et al., 2024). In parallel, LLM-based value identification has emerged as a lightweight alternative for text-only settings (Zhu et al., 2025).
2.2 Benchmarks and shared tasks for value detection
The field has converged on shared benchmarks to enable controlled comparisons and expose realistic difficulty factors such as multi-label outputs, class imbalance, and cross-domain variation. For argumentation, Kiesel et al. (2022) introduce a value-annotated benchmark for identifying human values behind arguments, and the Touché/ValueEval line of tasks systematizes evaluation and reporting (Mirzakhmedova et al., 2024; Kiesel et al., 2023). In Touché 2024, the Human Value Detection task frames value detection as sentence-level prediction under operational constraints typical of applied NLP pipelines (The ValuesML Team, 2024; Touché, 2024).
Within this context, recent systems emphasize cascaded decision processes and threshold control under label sparsity. The best-performing English system reported for Touché/CLEF 2024 uses a cascade to structure decisions and reduce spurious positives (Yeste et al., 2024). Yeste and Rosso (2026) study sentence-level value detection with moral presence gating and compute-frugal transformer ensembles, providing a strong baseline that we extend by focusing on Schwartz HO categories and their use as hierarchical structure.
Beyond shared tasks, recent resource suites aim for broader coverage across frameworks and domains; for example, MoVa aggregates multiple labeled datasets and benchmarks across moral/value theories to enable more generalizable evaluation (Chen et al., 2025).
2.3 Hierarchical structure and multi-label learning for value prediction
Our core question—whether HO structure helps basic-value prediction—connects to two mature ML literatures: multi-label learning and hierarchical classification. Multi-label classification is difficult when labels are non-mutually exclusive, skewed, and supported by limited positive evidence (Zhang and Zhou, 2014). Hierarchical classification studies how label taxonomies can share statistical strength and impose structure. These insights motivate our HO-aware variants: HO labels may reduce effective sparsity, but rigid enforcement can increase error propagation when parent predictions are uncertain (Silla and Freitas, 2011).
A related theme is the role of context when a single sentence provides limited evidence. Hierarchical document models (e.g., sentence-then-document encoders) improve classification by modeling multi-granular context (Yang et al., 2016). This suggests HO structure may be most effective as guidance combined with careful control of context, rather than as a brittle hard constraint.
2.4 Calibration, thresholding, and ensemble robustness under imbalance
Because value detection is multi-label and imbalanced, performance is often dominated by decision rules that map scores to binary labels. Classical calibration work (e.g., Platt scaling; Platt (1999)) and later studies on probability reliability show how miscalibration can distort precision/recall trade-offs, especially with per-label thresholds (Silva Filho et al., 2023). This motivates our emphasis on threshold tuning as a compute-frugal but high-leverage component.
Ensembling is another long-standing route to robustness under limited data and noisy supervision (Dietterich, 2000; Rokach, 2010). In value detection, different models can capture complementary cues, so small ensembles often deliver consistent gains without increasing single-model capacity. We build on this principle and the compute-frugal ensemble methodology in Yeste and Rosso (2026) to test whether HO-aware modeling improves beyond calibration and modest diversity.
2.5 Transformers and instruction-tuned LLMs for moral/value classification
Modern value detection systems typically rely on transformer encoders fine-tuned for classification (Devlin et al., 2019; He et al., 2021). Instruction-tuned LLMs have popularized prompt-based classification as a lightweight alternative that avoids task-specific architectural changes (Zhao et al., 2021; Wei et al., 2022; Liu et al., 2023). Work comparing prompting and supervised adaptation for human values motivates treating prompting vs. fine-tuning as an explicit design choice (Sun, 2024). For sentence-level, sparse multi-label settings, prompt-based LLMs still face calibration and recall challenges, often with less control over score distributions. Parameter-efficient methods such as QLoRA offer a middle ground between pure prompting and full fine-tuning (Hu et al., 2022; Dettmers et al., 2023).
Related work also examines how LLMs represent and align with human values, providing datasets and measurement frameworks. Examples include mapping LLM outputs to Schwartz value dimensions and constructing value-labeled data (Yao et al., 2024), measuring contextual alignment between human and LLM values across scenarios (Shen et al., 2025), and psychometric-style measurement of human/LLM values from text (Ye et al., 2025). Recent analyses probe value consistency and cultural alignment in LLMs (Rozen et al., 2025; Segerer, 2025; Biedma et al., 2024).
Taken together, these threads motivate the design space evaluated in this paper. We adopt the shared-task framing and compute-frugal discipline established in The ValuesML Team (2024); Touché (2024); Yeste and Rosso (2026), and we focus the comparison on how HO structure is injected (conditioning vs. hard gating/cascades) relative to strong, practically motivated baselines based on calibration and small ensembles. This sets up the methodological choices introduced next (Section 3).
3 Methodology
3.1 Problem formulation and label spaces
We study sentence-level human value detection under Schwartz’s refined theory (Schwartz, 2012). Each sentence may express none, one, or multiple values. Let be a sentence and the binary vector over the 19 basic values.111The benchmark provides attained and constrained annotations per value; we collapse them into a single expressed value signal (Section 3.2).
HO categories
To test whether coarser abstractions help, we deterministically derive eight HO binary labels from the 19 values following Schwartz (2012): Openness to Change, Conservation, Personal Focus, Social Focus, Self-Enhancement, Self-Transcendence, Growth, and Self-Protection. Let be the set of HO categories and the basic values grouped under . We define:
This yields . The value-to-HO mapping is fixed by theory and reported in A.
Presence label
We also consider a binary Presence gate that flags whether any value is expressed:
Bipolar evaluation slices
In addition to performance over all eight HO labels, we analyze the four canonical bipolar pairs to expose asymmetries in learnability and error propagation: (i) Openness to Change vs. Conservation, (ii) Self-Enhancement vs. Self-Transcendence, (iii) Personal Focus vs. Social Focus, and (iv) Growth vs. Self-Protection (Schwartz, 2012). For each pair, we report Macro- averaged over the two poles.
Figure 2 summarizes the relationship between the sentence input and the three label spaces.
3.2 Dataset and preprocessing
We use the official train/validation/test split (The ValuesML Team, 2024; Touché, 2024) at sentence level. To keep modeling choices unified and compute controlled, we use the benchmark’s English version (machine-translated sentences provided by the benchmark). This yields sentences: train, validation, and test. The split is at the text level; all sentences inherit their source text’s split.
Detailed label prevalence statistics (basic values and derived HO categories) for each split are reported in B.
Label construction
For each of the 19 values, the benchmark provides attained and constrained signals with values in , where denotes unclear. We binarize by treating any non-zero annotation as evidence of expression and collapse attained/constrained into a single label: that is, a value is marked as expressed if either its attained or constrained signal is non-zero.
3.3 Model families
We study three compute-frugal model families under a common protocol: (i) supervised transformer encoders, (ii) instruction-tuned LLMs used via prompting (zero-/few-shot), and (iii) parameter-efficient LLM fine-tuning (QLoRA). Predictions are evaluated as multi-label decisions over either the 19 values or the eight HO categories.
3.3.1 Direct multi-label prediction (supervised encoder)
The Direct approach follows Yeste and Rosso (2026). Given the pooled sentence representation , we apply a linear layer to produce logits and probabilities . Training minimizes standard multi-label binary cross-entropy: we minimize standard multi-label binary cross-entropy over labels. We do not use class weights, matching Yeste and Rosso (2026).
3.3.2 CategoryValues hierarchy (HO gating)
To test whether HO structure acts as an inductive bias, we implement a two-stage pipeline:
a)
Category stage: predict over the eight HO categories.
b)
Values stage: predict the basic values, but condition decisions on the category predictions.
We implement conditioning as a hard mask that constrains which basic values can be positive. If value belongs to HO category , we set:
where is the tuned threshold for category (Section 3.6). This makes a value permissive only when its parent HO category is predicted present.
3.3.3 PresenceCategoryValues cascade
We test a three-stage cascade where Presence acts as a first gate. The Presence formulation follows Yeste and Rosso (2026); the novelty here is the PresenceHOValues cascade:
a)
Presence stage: predict .
b)
Category stage: if Presence is positive, predict ; otherwise, output zeros for all HO labels.
c)
Values stage: apply the category-conditioned procedure from Eq. (3).
This cascade can improve precision by suppressing spurious positives on non-value sentences, but it can compound errors across stages. We quantify overall and per-pair effects in Section 3.8.
3.3.4 Instruction-tuned LLM baselines (definition prompting)
We benchmark instruction-tuned open LLMs that fit on a single 8 GB GPU (Llama 3.1 8B, Ministral 8B 2410, Qwen 2.5 7B, Gemma 2 9B). We use the definition-style prompt (best in Yeste and Rosso (2026)), presenting one-line definitions for the 19 values from Schwartz (2012). The model returns only a JSON array of applicable value names.
We evaluate zero-shot and few-shot prompting with in-context examples. Few-shot templates prepend exemplars in the same schema and include at least one null exemplar (empty array) when . We use greedy decoding with max_new_tokens=200.
Post-processing and label mapping
LLM outputs are parsed as JSON arrays and mapped to a multi-hot vector by exact string matching. Invalid generations (non-JSON or out-of-vocabulary labels) are treated as empty predictions. For HO evaluation, we derive from using Eq. (1) to keep model families comparable.
3.3.5 QLoRA fine-tuning (parameter-efficient LLM adaptation)
We additionally evaluate supervised QLoRA (Dettmers et al., 2023), following Yeste and Rosso (2026). Based on validation screening, we fine-tune Gemma 2 9B as the backbone. We train only low-rank adapters (base model frozen) with learning rate and save only adapter weights.
We train two QLoRA variants: (i) QLoRA direct, predicting the 19 values with rank and , three epochs, gradient accumulation 8, cosine schedule, and max length 512; and (ii) QLoRA hier, with rank , , three epochs, and max length 256. In both cases, LoRA targets the attention projection modules .
For QLoRA models that output probabilities, thresholds are tuned on validation under the same protocol as supervised encoders (Section 3.6) and then frozen for test.
Figure 3 summarizes the three main variants: a single-stage Direct predictor, a two-stage CategoryValues hierarchy with HO hard masks, and a three-stage PresenceCategoryValues cascade where a Presence gate filters sentences before HO and value prediction. All variants share the same encoder family and differ only in decision structure.
3.4 Compute-frugal auxiliary signals
Beyond the plain transformer baseline, we evaluate compute-frugal add-ons that are inexpensive relative to the encoder forward pass and fit within an 8 GB GPU budget:
Short local context. We concatenate up to the two previous sentences from the same source text to the current sentence (order preserved, separated by the model’s separator token, e.g., [SEP]), then truncate to length 512. We also attach a vector encoding their value labels (gold at training; model predictions at validation/test in an auto-regressive manner). This vector is projected to a low-dimensional embedding and concatenated with the text representation to provide short-range discourse cues.
Lexicon-derived features. We build sentence vectors from psycholinguistic, affective, moral, and value resources: LIWC-22 (Boyd et al., 2022), eMFD (Hopp et al., 2021), the Schwartz value lexicon in ValuesML (Kiesel et al., 2023), and affective lexica including NRC VAD (Warriner et al., 2013; Mohammad, 2018), NRC EmoLex (Mohammad and Turney, 2013), NRC Emotion Intensity (Mohammad and Kiritchenko, 2018), and WorryWords (Mohammad, 2024). We aggregate token signals into sentence statistics (counts, relative frequencies, averaged intensities) and standardize using training-set statistics.
When auxiliary features are enabled (supervised encoders), we concatenate them to the pooled transformer representation before classification. Unless stated otherwise, features are computed from the same input sentence (plus optional short local context).
3.5 Training protocol and compute parity
Unless otherwise stated, we follow the supervised protocol of Yeste and Rosso (2026). All transformer models fine-tune microsoft/deberta-base (He et al., 2021) with a linear multi-label head. Inputs are tokenized and truncated/padded to length 512. We optimize with AdamW (Loshchilov and Hutter, 2019) and a linear schedule with warmup, using batch size 4, gradient accumulation 4 (effective batch size 16), learning rate , weight decay 0.15, and up to 10 epochs with early stopping on validation Macro- (patience 4). Dropout is 0.1.
To keep comparisons fair and compute-frugal, we fix the encoder backbone and max sequence length across supervised variants, and select model variants and thresholds on validation only. All runs fit within a single 8 GB GPU. For LLM prompting, we use quantized decoding where applicable; for QLoRA we fine-tune adapter weights only.
To keep the study compute-frugal while covering many strategies (direct prediction, hard gates/cascades, auxiliary signals, prompting, QLoRA, ensembles), we fix a single random seed (as in Yeste and Rosso (2026)) for supervised runs. This prioritizes breadth under a fixed budget and keeps differences tied to modeling choices rather than to random initialization. Instead of multiple seeds, we emphasize paired evaluation: nonparametric sentence-level bootstrap uncertainty for Macro- and per-label paired tests (McNemar with FDR correction). We interpret differences below roughly 1–2 Macro- points conservatively and focus on effects supported by paired tests.
3.6 Threshold calibration
We map predicted probabilities to binary decisions using thresholds tuned on validation and then frozen for test. Let be the predicted probability for label and .
For supervised encoders (and QLoRA models that output probabilities), we use (i) a fixed global threshold or (ii) label-wise thresholds from a constrained grid search over . For each label , we maximize validation recall subject to precision (ties broken by higher recall). For hierarchical models, thresholds are tuned in stage-aware order (Presence first, then HO, then values). Prompted LLMs output discrete label sets and do not require threshold calibration.
3.7 Ensembling
To test whether small diversity gains yield robust improvements, we evaluate simple ensembles over a pool of trained models:
Hard voting: average binary decisions.
Soft voting: average predicted probabilities, then threshold (for families that output probabilities).
Weighted voting: probability averages weighted by validation Macro- (probability-outputting families only).
We build ensembles via forward selection: start from the best single model on validation, then add a candidate only if it improves validation performance and the one-sided bootstrap lower 95% bound for versus the current ensemble is and at least in relative terms; ties go to the smaller ensemble. For prompted LLMs (discrete outputs), we use hard voting; for supervised encoders (and QLoRA), we also use soft and weighted voting.
3.8 Evaluation metrics and statistical testing
Primary metric
We report macro-averaged over the label set. We compute per label across all sentences, then average across labels. For bipolar slices (Section 3.1), we compute Macro- for each pole and average the two poles.
End-task evaluation
End-task (end-to-end) evaluation is the main metric: we compute Macro- on the full evaluation split using the final system outputs for the target label space (e.g., the 19 values), where negative gate decisions force downstream predictions to zero (Eq. (3)). This captures both downstream quality and error propagation from upstream gating.
Uncertainty and significance
To assess robustness, we use nonparametric bootstrap resampling over sentences (Efron, 1992). We draw samples with replacement, recompute Macro-, and estimate (i) a one-sided 95% lower bound for and (ii) a one-sided empirical -value.
Per-label paired tests
3.9 Reproducibility
We log all configurations (preprocessing, model variants, thresholds, ensemble membership) and preserve prediction files for all runs. We also release the value-to-HO mapping and scripts for preprocessing and evaluation to enable exact replication on the benchmark splits (Touché, 2024). For LLM experiments, we release the prompts/templates and any adapter weights (where redistribution is permitted), along with deterministic decoding settings and post-processing scripts.
Figure 4 summarizes the pipeline. We start from the English ValueEval’24/ValuesML release, construct basic-value, HO, and Presence labels, and use the official train/validation/test splits. We then train or evaluate three model families under an 8 GB GPU budget (supervised encoders, prompted instruction-tuned LLMs, QLoRA-adapted LLMs), calibrate thresholds on validation, and select champion models. Finally, we form small ensembles and evaluate all systems with macro- and paired significance tests (bootstrap and McNemar) on both HO and basic-value slices.
4 Results
We organize results around the research questions from Section 1: we first address HO category learnability (RQ1) and compute-frugal upgrades (RQ4), then evaluate hierarchical mechanisms for downstream value prediction (RQ2–RQ3), and finally benchmark small instruction-tuned LLMs under the same budget (RQ5).
For completeness, Section S2 reports the full validation/test tables for all higher-order category experiments and ablations. Here we focus on the main trends.
4.1 Higher-order categories are learnable, but not equally so
HO categories are learnable with compact supervised encoders, but difficulty varies by pair. The easiest pair is Growth vs. Self-Protection (test Macro-, Table S3). Self-Transcendence vs. Self-Enhancement is moderate (best Macro-, Table S9), while Openness vs. Conservation is hardest (best Macro-, Table S7). These patterns track label prevalence: Openness is rare (about 8% vs. about 20% for Conservation, Table 2), which likely drives both lower Macro- and strong pole asymmetry (Openness Conservation in Table 1).
Table 1 summarizes these differences: Growth vs. Self-Protection is most learnable (Macro-), Self-Transcendence vs. Self-Enhancement is moderate (best Macro-), and Openness vs. Conservation remains hardest (best Macro-), with persistent asymmetry (Conservation Openness). The pattern suggests that “constraint/tradition” cues are captured more reliably than “novelty/autonomy” cues at sentence level. We report single-run results and interpret differences under 1–2 Macro- points cautiously.
| HO pair | Baseline | Best model | Best tuned |
| Fixed/Tuned | (if baseline) | /pole A/pole B | |
| Growth vs. Self-Protection | 0.58/0.58 | Baseline | 0.58/0.54/0.62 |
| Social Focus vs. Personal Focus | 0.41/0.43 | NER / WorryWords / LIWC15 | 0.57/0.59/0.54 |
| Openness to Change vs. Conservation | 0.42/0.38 | LIWC22 + Ling. Feat | 0.42/0.34/0.50 |
| Self-Transcendence vs. Self-Enhancement | 0.48/0.50 | WorryWords / MFD-20 | 0.51/0.53/0.48 |
A second asymmetry appears in per-class : for harder pairs, one pole is much more learnable (e.g., Conservation consistently outperforms Openness, Table S7). This suggests that models capture “normative constraint” language (rules, tradition, order) more reliably than implicit “novelty/autonomy” cues.
4.2 Cheap knobs matter: threshold calibration is consistently helpful, except when it overfits
Table 1 shows the effect of fixed vs. tuned thresholds: tuning strongly helps Social Focus vs. Personal Focus (0.41 0.57) and yields a smaller but consistent gain for Self-Transcendence vs. Self-Enhancement (0.48 0.51), while it can overfit under severe imbalance for Openness vs. Conservation (0.417 0.38 for the baseline).
Threshold calibration is a low-cost lever with frequent gains. For Social Focus vs. Personal Focus, tuned thresholds consistently improve Macro-, reaching on test (e.g., NER/WorryWords/LIWC15, Table S5). The fixed-0.5 baseline underperforms (0.41), suggesting sensitivity to calibration and/or distribution shift; tuned thresholds recover performance (Table 1). For Self-Transcendence vs. Self-Enhancement, tuning improves from to , with best variants at (Table S9). In contrast, for Openness vs. Conservation, tuned thresholds can reduce Macro- (Table S7), consistent with per-label overfitting under severe imbalance (e.g., extreme thresholds for Openness).
Overall, label-wise thresholds are usually worth it in compute-frugal regimes, but highly imbalanced labels benefit from conservative tuning rules (or calibration methods that regularize thresholds toward a global prior).
4.3 Lightweight auxiliary signals rarely move the needle, but can stabilize certain pairs
Most feature add-ons (lexica, topic features, short context) yield small or inconsistent improvements and rarely beat a well-tuned baseline. Still, a few trends emerge.
Table 1 shows that auxiliary signals change top rankings for only a subset of pairs: Social Focus vs. Personal Focus benefits from weakly supervised cues (NER and lexical resources reach with tuned thresholds), and Self-Transcendence vs. Self-Enhancement shows modest gains (best ). In contrast, Growth vs. Self-Protection is near its ceiling, and Openness vs. Conservation remains difficult even with added features.
For Social vs. Personal, weakly supervised cues help at test time (Table S5): NER and affective/moral lexica (e.g., WorryWords, LIWC15) reach Macro-. This likely reflects explicit references to social groups, institutions, or interpersonal relations that lexical and NER cues partially capture. For Self-Transcendence vs. Self-Enhancement, lexical signals again help modestly (Macro-, Table S9), suggesting a small benefit from prosocial vs. status/power cues.
At the same time, ablations show that auxiliary signals can hurt sharply if misaligned or noisy. This may reflect configuration issues (e.g., feature scaling or dimensional mismatch), but the takeaway is practical: cheap features are only cheap if they are robustly engineered; otherwise they can dominate the classifier head and destabilize training.
4.4 Presence gating boosts in-gate validation scores but does not robustly improve end-task performance
Presence gating (Eq. (2)) produces a large jump in in-gate validation Macro- when evaluation is restricted to sentences predicted to contain a value (Section S3). For example, Growth vs. Self-Protection rises from about to , and Social vs. Personal from about to . This is expected: removing many easy negatives changes the operating point and reduces sparsity.
Table 2 shows the same pattern across all four HO pairs: restricting evaluation to gate-passing sentences increases validation Macro- by +0.14 to +0.16, but these gains do not translate into consistent improvements on the original test distribution.
| HO pair | Val Macro- | Test Cascading Macro- |
| Direct / Presence gate | Direct / Presence-gated | |
| Growth vs. Self-Protection | 0.58 / 0.77 | 0.58 / 0.58 |
| Social Focus vs. Personal Focus | 0.54 / 0.74 | 0.57 / 0.56 |
| Openness to Change vs. Conservation | 0.42 / 0.58 | 0.42 / 0.43 |
| Self-Transcendence vs. Self-Enhancement | 0.45 / 0.59 | 0.51 / 0.50 |
However, on the original test distribution, Presence gating does not yield consistent improvements. The best gated Macro- matches the direct baseline for Growth/Self-Protection (0.58), slightly underperforms for Social/Personal (0.56 vs. 0.57) and Self-Transcendence/Self-Enhancement (0.50 vs. 0.51), and yields only a marginal gain for Openness/Conservation (0.43 vs. 0.42) (Table 2).
The most plausible explanation is error compounding: Presence false negatives suppress downstream positives (recall loss), while Presence false positives admit sentences that still look negative to the HO classifier (precision loss). The net effect is that hard gating trades recall for precision in a way that is hard to tune globally. This sensitivity is visible empirically: changing the gate threshold (e.g., 0.50 vs. 0.10) often leaves test Macro- unchanged or slightly worse (Tables S11 and S13–S17). This motivates softer alternatives rather than a binary filter.
Paired tests corroborate these trends: in several HO slices the best tuned Direct systems are significantly stronger than Presence-based alternatives on test (Section S7), indicating that PresenceCategory cascades do not provide a robust out-of-sample advantage under hard gating.
4.5 HOvalues hard gating does not translate into reliable gains on fine-grained detection
We next test whether HO predictions help fine-grained value detection by constraining the value space. Table 3 summarizes the best test Macro- with hard HOvalues gates across all HO categories (details in Section S4). Overall, hard HO gating via a binary mask is not a “free win.” Even the best gated configurations remain modest and do not consistently outperform tuned Direct baselines (Section S7).
| HO gate | #values | Best test Macro- | Best configuration (value model / gate) |
| Growth | 11 | 0.271 | Baseline@0.30 / Baseline@0.50 |
| Self-Protection | 10 | 0.306 | Baseline@0.15 / Baseline@0.50 |
| Social Focus | 10 | 0.326 | Lex - MJD@0.25 / TD - NMF@0.50 |
| Personal Focus | 9 | 0.287 | Lex - MJD@0.70 / NER@0.19 |
| Openness | 4 | 0.235 | NER + TD - LDA@0.50 / Lex - LIWC 22 + Ling Feat@0.50 |
| Conservation | 7 | 0.297 | Lex - Schwartz@0.50 / Lex - LIWC 22 + Ling Feat@0.10 |
| Self-Transcendence | 6 | 0.307 | Baseline@0.25 / Lex - MFD-20@0.50 |
| Self-Enhancement | 5 | 0.320 | Lex - Schwartz@0.55 / Lex - WorryWords@0.23 |
Under the Growth gate, the best test Macro- for Growth values is about ; under Self-Protection it is about . Section S7 compares the best HO-gated systems against tuned Direct baselines within each HO slice using paired bootstrap tests. The pattern is consistent: hard HO gating does not yield reliable improvements and is sometimes significantly worse, consistent with recall suppression when parent categories are missed. In a sentence-level setting—where signals are short, implicit, and often multi-valued—this loss is hard to recover.
In short, HO structure is informative, but forcing predictions to respect the hierarchy via a binary mask is often too rigid for noisy sentence-level supervision.
4.6 Where small instruction-tuned LLMs fit under the same budget
Section S4 benchmarks instruction-tuned 10B LLMs (prompted and QLoRA-adapted) on the same HO-restricted slices. Table 4 reports the best test Macro- for Gemma-2-9B-it under (i) prompting only, (ii) prompting plus a lightweight SBERT gate, and (iii) QLoRA adaptation, and indicates which upgrades are supported by bootstrap tests (Section S7). Overall, even the best LLM variants remain well below the supervised DeBERTa-based champions on test (e.g., Growth: vs. ; Self-Protection: vs. ; Social Focus: vs. ; Personal Focus: vs. ; see Tables 4, 5).
| HO slice | Prompted | +gate | QLoRA | Ensemble | Sig. |
| Growth | 0.191 | 0.194 | 0.132 | 0.201 | FS+, G+, Q0, E+, X– |
| Self-Protection | 0.220 | 0.212 | 0.254 | 0.254 | FS+, G0, Q+, E–, X+ |
| Social Focus | 0.230 | 0.234 | 0.206 | 0.267 | FS+, G+, Q0, E+, X– |
| Personal Focus | 0.198 | 0.195 | 0.184 | 0.198 | FS+, G0, Q0, E–, X+ |
| Openness | 0.093 | 0.113 | 0.123 | 0.123 | FS+, G+, Q0, E–, X0 |
| Conservation | 0.230 | 0.227 | 0.258 | 0.258 | FS+, G+, Q+, E–, X+ |
| Self-Transcendence | 0.213 | 0.221 | 0.124 | 0.221 | FS+, G+, Q0, E0, X– |
| Self-Enhancement | 0.221 | 0.222 | 0.227 | 0.227 | FS+, G–, Q–, E–, X– |
Bootstrap comparisons in Section S7 show that (i) few-shot prompting significantly improves over zero-shot in multiple slices (FS+; e.g., Growth, Social Focus, Self-Protection, Personal Focus), and (ii) adding an SBERT-style gate can yield additional gains in some settings (G+; e.g., Growth and Social Focus), but not reliably in others (e.g., Personal Focus, Self-Protection). QLoRA adaptation is mixed: it improves in Self-Protection and Conservation, but degrades in Growth and Social Focus, suggesting sensitivity to sparsity and slice-specific shifts. Where reported, within-family LLM ensembling can also help (E+; Growth, Social Focus).
LLMs can still be useful as a diversity source in cross-family ensembles. For Self-Protection and Personal Focus, combining the best transformer with the best LLM yields a further significant improvement (X+), indicating complementary error patterns despite weaker standalone performance.
4.7 Simple ensembling is the most reliable compute-frugal gain
Across HO slices, the strongest and most repeatable improvements come from small, low-cost ensembles rather than hard hierarchical masks. Table 5 summarizes test Macro- for the transformer champions ensemble (soft voting) and the corresponding bootstrap outcomes (Sections S6 and S7). Overall, ensembling yields consistent point gains, with the clearest improvements in Growth, Self-Protection, and Personal Focus.
| HO slice | Single | Tr ens.; Sig. | Tr+LLM; Sig. |
| Growth | 0.286 | 0.303; + | –; – |
| Self-Protection | 0.321 | 0.342; + | 0.353; + |
| Social Focus | 0.330 | 0.345; 0 | –; – |
| Personal Focus | 0.305 | 0.317; + | 0.326; + |
| Openness to Change | 0.240 | 0.241; 0 | 0.249; 0 |
| Conservation | 0.301 | 0.313; 0 | 0.334; + |
| Self-Transcendence | 0.302 | 0.319; 0 | –; – |
| Self-Enhancement | 0.338 | 0.337; – | 0.349; 0 |
Section S7 confirms that these ensemble gains are not just noise. For example, in Growth, moving from the tuned Direct baseline to the transformer ensemble increases Macro- from to and is significant. Self-Protection and Personal Focus show similar significant lifts over the best single model in the paired comparison.
Not all slices benefit equally: for Social Focus, the ensemble improves the point estimate but does not meet the one-sided bootstrap criterion. Similar non-significant (but positive) gains appear in Openness, Conservation, and Self-Transcendence (Table 5).
4.8 Which improvements are statistically robust under fixed compute
Section S7 evaluates paired differences with a one-sided bootstrap test on Macro- and per-label McNemar tests with Benjamini–Hochberg correction. Table 6 condenses the main fixed-compute robustness results across all slices. Three consistent patterns emerge.
| HO Slice | Direct vs gated; Sig. | Ensemble; Sig. | Hybrid; Sig. |
| Growth (HO) | 0.003; + | 0.002; + | – |
| Self-Protection (HO) | 0.002; + | 0.006; + | 0.007; + |
| Social Focus (HO) | -0.002; 0 | -0.001; 0 | – |
| Personal Focus (HO) | 0.011; + | 0.003; + | 0.005; + |
| Openness (HO) | -0.003; 0 | -0.016; 0 | -0.010; 0 |
| Conservation (HO) | -0.009; 0 | -0.006; 0 | 0.011; + |
| Self-Transcendence (HO) | -0.016; 0 | -0.007; 0 | – |
| Self-Enhancement (HO) | 0.010; + | -0.014; 0 | -0.003; 0 |
First, threshold tuning is a statistically reliable improvement: in every slice with reported tests, tuned thresholds significantly outperform fixed for Direct models (see “Thr. tuning” in Table 6). McNemar analyses show these gains concentrate in subsets of labels (e.g., Universalism/Beneficence in Self-Transcendence; Security/Conformity/Tradition in Conservation; several sparse HO labels), rather than uniformly.
Second, hard hierarchical gating is not a reliable downstream win. For HO slices, the tuned Direct champion significantly outperforms the HO-gated champion in three of four cases (Growth, Self-Protection, Personal Focus); Social Focus shows no significant difference. For HO categories, none of the reported gated champions beats the Direct champion (Openness, Conservation, Self-Transcendence), consistent with error compounding under hard masks.
Third, ensembles provide the most consistent significant gains beyond threshold tuning, mainly in HO slices. Transformer soft-voting ensembles yield significant improvements in Growth, Self-Protection, and Personal Focus, while Social Focus (HO) and all reported HO categories show non-significant uplift. Hybrid ensembles can yield additional gains in some cases (Self-Protection, Personal Focus, Conservation), but not universally (e.g., Openness).
5 Discussion
Taken together, the experiments point to five practical outcomes.
(1) HO abstractions are learnable, but they are not uniformly reliable
Pairs with higher prevalence and stronger lexical regularities (e.g., Growth/Self-Protection) are easier; rare or diffuse categories (Openness) remain difficult even with tuning. For example, Growth/Self-Protection reaches Macro- , while Openness/Conservation peaks around with persistent pole asymmetry (Conservation Openness).
(2) Calibration and small ensembles are safer bets than hard hierarchies
Threshold tuning yields small but frequent gains, and forward-selected soft-voting ensembles provide the most consistent significant improvements (Section S7), while most feature add-ons are marginal or unstable. In HO detection, tuning ranges from modest gains (e.g., 0.48 0.51) to large calibration-sensitive jumps (e.g., Social/Personal 0.41 0.57); ensembles yield smaller but more reliable lifts (e.g., Growth 0.286 0.303).
(3) Presence gating and HO gating improve conditional performance but not end-task performance
Large validation gains under gating are largely an artifact of evaluating a simplified subproblem (value-present sentences) and do not carry over on the full test distribution. Presence gating inflates in-gate validation Macro- by roughly +0.14 to +0.16, but test improvements are negligible or negative in most slices. A unifying explanation is error compounding: parent false negatives suppress child recall, parent false positives admit hard negatives, and imbalance makes threshold search volatile for rare poles (e.g., Openness).
(4) Hard gates do not yield reliable end-to-end gains
We directly test hard hierarchical mechanisms (Presence gating and HOvalues masking). Across HO pairs and value slices, these hard constraints do not yield reliable end-to-end gains and are sometimes significantly worse than tuned Direct models, despite strong conditional scores. This is consistent with error propagation: uncertain parent decisions become binary filters that suppress true positives and hurt recall for sparse labels. While we do not evaluate soft hierarchical conditioning here, the results motivate treating HO structure as an uncertainty-preserving inductive bias (e.g., probabilistic conditioning or auxiliary HO objectives) rather than a strict routing rule, and exploring broader context to reduce ambiguity.
(5) Small LLMs are not competitive alone, but can add useful diversity
Under the same budget, prompted and QLoRA-adapted 10B LLMs underperform supervised encoders in absolute Macro-, although few-shot prompting helps. Their main practical benefit is as complementary signals in cross-family ensembles for some slices (Section S7). For instance, Gemma-2-9B-it remains below transformer champions (e.g., Growth vs ), but can still improve cross-family ensembles in selected slices.
Beyond this benchmark, our results show that how domain knowledge is injected matters as much as which knowledge is used. Enforcing the HO taxonomy as a hard constraint may raise precision in a restricted space but can reduce end-task recall through error propagation. From a system-design standpoint, psychologically grounded taxonomies such as Schwartz values are best leveraged as regularizers and priors rather than strict filters when predictions are noisy or labels overlap. Overall, enforcing hierarchy via hard gating is brittle in sentence-level, imbalanced, multi-valued settings.
5.1 Limitations and threats to validity
Our findings have several limitations. First, we report single-run results, so differences of 1–2 Macro- points may be unstable. Second, the sentence-level setting limits signal; hierarchy may help more with broader discourse context. Annotation noise and multi-label overlap can also conflict with strict parent–child constraints. Third, calibration can overfit under severe imbalance, especially for rare categories (e.g., Openness). Finally, conclusions are scoped to this benchmark and a compute-frugal regime; results may shift with more data, new domains, alternative schemes, or larger models.
5.2 Answers to the research questions
RQ1 (Are HO values learnable from single sentences?). Yes—HO categories are learnable with compact supervised encoders, but learnability varies widely across pairs; rare/diffuse categories (e.g., Openness) remain challenging under fixed compute (Section 4.1).
RQ2 (Do HO gates help downstream basic-value prediction?). Under hard masking (CategoryValues), HO gating does not reliably improve out-of-sample Macro- and can be significantly worse than tuned Direct models (Section S7), consistent with error compounding. This negative result is specific to hard masking and does not rule out gains from soft HO integration, which we do not evaluate here.
RQ3 (Does PresenceCategory outperform Category-only?). With hard Presence-gated cascades, Presence improves conditional performance but the full pipeline does not consistently beat tuned Direct baselines on the test distribution (Section 4.4). Gains are not robust across slices (Section S7). This negative result is specific to binary gating and does not preclude improvements from learned or soft gates or end-to-end training.
RQ4 (Which low-cost knobs move the needle?). Threshold calibration is the most consistently significant improvement, and simple soft-voting ensembles provide additional gains in several HO slices (Sections 4.7–4.8). Lexica/topic/context features are unstable: they can help specific slices but are not the main drivers under fixed compute.
RQ5 (Where do small LLMs fit?). Prompted 10B LLMs benefit from few-shot prompting and sometimes from lightweight semantic gates, but they lag behind supervised DeBERTa-based models under the same budget (Section 4.6). Their practical value is mainly as complementary signals in cross-family ensembles, which can yield significant improvements in some slices (Section S7).
6 Conclusions and future work
This paper examined whether higher-order (HO) value abstractions improve fine-grained, sentence-level human value detection under a compute-frugal regime. Across encoder-based systems and compact instruction-tuned LLMs, we tested HO-aware strategies spanning direct HO prediction, HOvalue inference via hard gating/cascades, and low-cost upgrades such as label-wise threshold calibration and small soft-voting ensembles.
Our results yield three conclusions. First, HO categories are learnable from single sentences, but learnability is uneven and correlates with prevalence and lexical regularities: pairs such as Growth/Self-Protection are easier than rare or diffuse categories such as Openness, which also shows persistent pole asymmetries. Second, the most reliable gains come from calibration and ensembling, not from handcrafted features or strict hierarchical inference. Label-wise threshold tuning helps often (but can overfit under severe imbalance), and small soft-voting ensembles provide the most consistent gains. Third, hard hierarchical mechanisms look strong under conditional evaluation but do not robustly improve the end task. Presence gating and HOvalues masking raise validation scores on gate-passing sentences, but gains largely disappear (or reverse) on the full distribution, consistent with error compounding and recall suppression under sparse supervision. HO structure is informative, but enforcing it as a hard constraint is often too brittle for noisy, multi-valued sentences.
Compact instruction-tuned LLMs are not competitive with supervised encoders in absolute Macro- under the same budget, but they can still help as complementary signals: in some slices, cross-family ensembles benefit from LLM diversity even when LLMs lag as standalone predictors. Overall, hierarchy is best used as an inductive bias that guides inference softly, rather than as a rigid routing rule, at least for sentence-level value detection with sparse multi-label targets.
Future work should replace brittle hard gates with approaches that preserve uncertainty and better match the annotation reality. A promising direction is joint hierarchical learning, where HO and fine-grained labels are predicted in a single model via multi-task objectives or structured decoders with shared representations. Closely related are soft HO priors that condition value predictions on HO probabilities (rather than masking), which can encourage coherence without catastrophic recall loss. Given instability under imbalance, more principled calibration for rare labels is also needed. Finally, because sentence-level inputs limit signal and values may be expressed across discourse, future work should systematically vary context (sentence, local window, document) to test which values benefit from broader evidence. Evaluations on additional domains and schemes will help clarify when value hierarchies help most and how to use them robustly in data-limited settings.
7 CRediT authorship contribution statement
Víctor Yeste: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Resources, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Project administration. Paolo Rosso: Supervision, Writing - Review & Editing.
8 Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the authors used ChatGPT from OpenAI to improve the readability and language of the manuscript. After using this tool/service, the authors reviewed and edited the content as needed. The authors take full responsibility for the content of the published article.
9 Data availability
This study uses the English, machine-translated ValueEval’24/ValuesML release (Mirzakhmedova et al., 2024; The ValuesML Team, 2024). The dataset is distributed under a Data Usage Agreement that allows research use but does not permit redistribution of the texts. Accordingly, we cannot share the sentences or any derivative file that contains original textual content. Researchers with appropriate access can obtain the official train/validation/test splits by registering and downloading the release from Zenodo (The ValuesML Team, 2024).
To maximize reproducibility without violating the license, we release the full experimental pipeline and all non-text artifacts needed to reproduce our results. Specifically, we release (i) the codebase for preprocessing, model training, evaluation, threshold calibration, and ensembling, with configuration files for architectures and hyperparameters; (ii) trained model artifacts, including fine-tuned DeBERTa checkpoints (direct predictors, gating components, feature-augmented variants) and QLoRA adapter weights for Gemma 2 9B; and (iii) inference outputs for every system (validation and test), including predicted probabilities, binarized decisions from selected thresholds, and the thresholds themselves (global and per-label where applicable). These resources are provided via GitHub222https://github.com/VictorMYeste/human-value-detection and Hugging Face333https://huggingface.co/papers/2601.14172.
All released files are keyed only by the dataset’s official identifiers (e.g., Text-ID and Sentence-ID) and contain no original sentences. This allows any researcher with licensed access to ValueEval’24 to reproduce our tables and figures and build further analyses on the same splits.
Appendix A Value-to-HO mapping
Table 1 reports the fixed mapping used in this paper between Schwartz’s 19 refined basic values and the eight HO categories. Note that in the refined theory some basic values contribute to more than one HO category (e.g., Hedonism, Achievement, Face, Humility), so overlaps across rows are expected.
| HO category | Basic values included |
| Growth | Humility; Benevolence: caring; Benevolence: dependability; Universalism: concern; Universalism: nature; Universalism: tolerance; Self-direction: thought; Self-direction: action; Stimulation; Hedonism; Achievement. |
| Self-Protection | Achievement; Power: dominance; Power: resources; Face; Security: personal; Security: societal; Tradition; Conformity: rules; Conformity: interpersonal; Humility. |
| Social Focus | Security: societal; Tradition; Conformity: rules; Conformity: interpersonal; Humility; Benevolence: caring; Benevolence: dependability; Universalism: concern; Universalism: nature; Universalism: tolerance. |
| Personal Focus | Self-direction: thought; Self-direction: action; Stimulation; Hedonism; Achievement; Power: dominance; Power: resources; Face; Security: personal. |
| Openness to Change | Self-direction: thought; Self-direction: action; Stimulation; Hedonism. |
| Self-Enhancement | Hedonism; Achievement; Power: dominance; Power: resources; Face. |
| Conservation | Face; Security: personal; Security: societal; Tradition; Conformity: rules; Conformity: interpersonal; Humility. |
| Self-Transcendence | Humility; Benevolence: caring; Benevolence: dependability; Universalism: concern; Universalism: nature; Universalism: tolerance. |
Appendix B Label prevalence across data splits
This appendix reports the prevalence of each label in the train/validation/test splits. Prevalence is computed at the sentence level as the percentage of sentences annotated with a given label. The Presence row corresponds to the percentage of sentences with at least one label.
| Label | Train | Validation | Test |
| Self-direction: thought | 1.29 | 1.15 | 1.17 |
| Self-direction: action | 3.61 | 3.26 | 3.51 |
| Stimulation | 2.62 | 2.82 | 2.55 |
| Hedonism | 0.86 | 0.67 | 0.86 |
| Achievement | 6.42 | 6.37 | 6.25 |
| Power: dominance | 4.63 | 4.40 | 4.33 |
| Power: resources | 5.00 | 4.86 | 5.53 |
| Face | 1.81 | 1.90 | 1.83 |
| Security: personal | 2.03 | 1.87 | 2.42 |
| Security: societal | 8.95 | 8.46 | 7.90 |
| Tradition | 1.20 | 1.84 | 1.35 |
| Conformity: rules | 6.10 | 6.41 | 6.25 |
| Conformity: interpersonal | 1.35 | 1.37 | 1.34 |
| Humility | 0.24 | 0.29 | 0.21 |
| Benevolence: caring | 2.29 | 2.29 | 2.22 |
| Benevolence: dependability | 1.94 | 1.93 | 1.98 |
| Universalism: concern | 4.97 | 4.50 | 5.04 |
| Universalism: nature | 2.05 | 2.57 | 2.01 |
| Universalism: tolerance | 1.07 | 0.81 | 1.17 |
| Presence | 51.53 | 50.99 | 50.81 |
| Dimension | Train | Validation | Test |
| Growth | 25.56 | 24.68 | 25.16 |
| Self-Protection | 35.20 | 35.41 | 34.82 |
| Social Focus | 28.19 | 28.48 | 27.50 |
| Personal Focus | 26.59 | 25.54 | 26.72 |
| Openness to Change | 8.20 | 7.72 | 7.90 |
| Conservation | 20.90 | 21.40 | 20.34 |
| Self-Transcendence | 12.16 | 11.93 | 12.27 |
| Self-Enhancement | 18.03 | 17.62 | 18.24 |